The radius and height of a cone are respectively r and h. If the height is multiplied by 9/2; by what fraction should the radius be multiplied so as to quadruple the cones volume?


1
Expert's answer
2021-11-01T07:18:15-0400
"V=\dfrac{1}{3}\pi r^2h""4V=\dfrac{1}{3}\pi (kr)^2(\dfrac{9}{2}h)""\dfrac{\dfrac{1}{3}\pi (kr)^2(\dfrac{9}{2}h)}{\dfrac{1}{3}\pi r^2h}=4""k^2=\dfrac{8}{9}""k=\dfrac{2\sqrt{2}}{3}""k=\dfrac{2\sqrt{2}}{3}"

The radius should be multiplied by "\dfrac{2\sqrt{2}}{3}."

"\dfrac{2\sqrt{2}}{3}."